# Planet Migration in Protoplanetary Disks with Rims

### v1

@@ 3.1 Jupiter Migration
There is a general trend that the planet would migrate outward when the rim is located at $\Delta r_{\rm dz\_out}>-\Delta r_{\rm dz\_in}$, while the planet would migrate inward if the rim location satisfies $\Delta r_{\rm dz\_out}<-\Delta r_{\rm dz\_in}$ (see Panel b in Fig. 1). There is nearly zero migration torque when the planet is located at the rim center where $\Delta r_{\rm dz\_out}\simeq-\Delta r_{\rm dz\_in}$, shown as the dotted line in the panel(c) of Fig. 1. This simply suggests the dead zone center ($r_{\rm mid}=(r_{\rm dz\_in}+r_{\rm dz\_out})/2$) as the critical boundary of the migration torque reversal. Spatial gradients in the color map reveal smooth variations in the migration torque in terms of both its amplitude and sign. The size of the symbol reveals the variance of the torque, which could be related to the vortex-induced torque variations, as we show below.
@@
3.2 Super-Earth Migration
@@
For super-Earths, we can apply semi-analytical calculations by the formula for type I migration (Section 2.2.2). The results for the comparison between theoretical and simulated torque values of the models enclosed by the green rectangle in Fig. 1(d) are illustrated in Fig. 4.
Figure 4: Theoretical calculation of the torque for super-Earth. Here, we set $M_{\rm p}=10M_{\oplus}$. We compute the theoretical torque values for the models enclosed by the green rectangle in Fig. 1(d). All these models have a dead zone width of $0.7r_{0}$, and the inner boundary of the dead zone $r_{\rm dz\_in}$ is shown on the horizontal axis of the figure. The blue diamond-dashed line represents the Lindblad torque, and the red square-dashed line denotes the corotation torque. The black dot-solid line shows the sum of the two obtained from analytical calculation, representing the total torque. The purple x-solid line corresponds to the total torque directly computed from Eq. 6.
@@ 3.3 Effect of Vortex
The formation of vortices in protoplanetary disks is primarily driven by instabilities that arise under specific disk conditions. Vortices can be generated through multiple mechanisms, including Rossby Wave Instability (RWI) at the edges of gaps carved by massive planets in low-viscosity disks, at the boundaries of accretionally inactive dead zones, via the influence of binary companions, or through baroclinic instability. The RWI (lovelace1999; lihui2000; lihui2001; Lyra2013) is a dominant mechanism responsible for generating the large-scale, dust-trapping vortices observed in transitional disks. The RWI is characterized as an “edge mode” instability, analogous to the Kelvin-Helmholtz instability, which operates at sharp pressure maxima within the disk. It efficiently converts excess shear into vorticity, leading to the formation of long-lived, coherent vortical structures (balmforth2001; koller2003; lihui2005; Li2020).
In this study, vortex formation results from the combined effects of Jupiter’s gap-opening process and the ring-shaped surface density bump. The gap carved by Jupiter produces an annular region of reduced surface density, while the annular surface density bump is an accumulation of mass. The relative position and interaction between these two features determine both the occurrence and strength of any resulting vortices. For vortices at dead zone edges is primarily driven by gas accumulates at this viscosity transition, to maintain a radially constant accretion rate (${\dot{M}}\sim\nu\Sigma_{g}$), leading to the formation of a local maximum in the gas pressure profile (a “bump”). When this pressure bump is sufficiently sharp (with a width $\Delta_{\rm DZ}\lesssim 2H$, where $H$ is the local scale height), it becomes unstable to the RWI, which non-linearly evolves into long-lived, anticyclonic vortices (Miranda2017). This vortex generation mechanism at the dead zone edge is self-sustaining because the ongoing accretion flow continuously replenishes the gas at the bump, allowing the RWI and the vortices to persist for thousands of orbits.
@@
3.4 Effect of Planet Accretion
@@
According to li2024, when the disk’s $\alpha$ parameter exceeds 0.003, accretion causes the planet to transition from inward to outward migration. In our setup, the dead zone has an $\alpha$ of $10^{-4}$, while the active region has an $\alpha$ of $10^{-2}$. The $\alpha$ in the active region exceeds the critical value required for accretion to exert an influence, whereas the $\alpha$ in the dead zone is insufficient for accretion to have a significant effect. Although accretion enhances the outward migration tendency of Jupiter, it is not sufficient to reverse the overall migration direction.
@@
3.5 The Cases with a Surface Density Gap
@@
Figure 7: The migration behavior of planets in a disk with a surface density gap. We inverted the $\alpha$ profile by swapping the values in the originally defined dead and active zones, while maintaining a smooth transition between them. The top row of panels (panel (a)–(c)) displays the migration map for Jupiter, whereas the bottom row (panel (d)–(f)) corresponds to super-Earths. For the upper column (left to right): the panel (a) employs $r_{\mathrm{dz\_in}}=0.8\,r_{0}$, $r_{\mathrm{dz\_out}}=1.4\,r_{0}$; the panel (b) uses $r_{\mathrm{dz\_in}}=0.7\,r_{0}$, $r_{\mathrm{dz\_out}}=1.3\,r_{0}$; and the panel (c) is specified as $r_{\mathrm{dz\_in}}=0.6\,r_{0}$, $r_{\mathrm{dz\_out}}=1.2\,r_{0}$. For the lower column (left to right): the panel (d) adopts $r_{\mathrm{dz\_in}}=1.0\,r_{0}$, $r_{\mathrm{dz\_out}}=1.6\,r_{0}$; the panel (e) has $r_{\mathrm{dz\_in}}=0.7\,r_{0}$, $r_{\mathrm{dz\_out}}=1.3\,r_{0}$; and the panel (f) corresponds to $r_{\mathrm{dz\_in}}=0.4\,r_{0}$, $r_{\mathrm{dz\_out}}=1.0\,r_{0}$. The selected configurations, indicating the relative positions between the surface density gap and the planet, are marked by diamond symbols in Fig. 1(c) and (d). The resulting migration directions, as shown in the left panel, are opposite to those shown in Fig. 1(b).
As shown in Fig. 7, panels (a)–(c) present the migration behavior of Jupiter in a disk with an annular surface density gap, while panels (d)–(f) show the corresponding results for super-Earths. The relative positions of the planets with respect to the surface density gap in these configurations are marked by diamond symbols in Fig. 1(c) and (d). For Jupiter, when located within the ring-shaped surface density gap, its migration direction is opposite to that observed in the annular surface density bump: with respect to the center of the surface density gap, planets migrate outward in the interior region and inward in the exterior region, causing Jupiter to converge toward the gap center. Moreover, gap opening by Jupiter further deepens and reinforces the pre-existing intrinsic surface density gap. For super-Earths, planets inside the ring-shaped surface density gap predominantly undergo inward migration. However, on or slightly outside the outer boundary of the surface density gap, outward migration occurs, whereas inward migration persists near the inner boundary. This implies that super-Earths are unlikely to remain trapped within the ring-shaped surface density gap, suggesting that intrinsic surface density gaps may not harbor super-Earths.
@@
4 Conclusion and Discussions
@@
Our calculations are scale-free; however, we can provide an example application to observational estimates. For instance, the cumulative torque in the top-left panel of Fig. 3 is approximately $T\sim 10^{-4}\Sigma_{0}r_{0}^{4}\Omega_{0}^{2}$. By adopting an orbital radius of $r_{0}=1\,\text{AU}=1.496\times 10^{13}\,\text{cm}$, a surface density of $\Sigma_{0}=1700\,\text{g/cm}^{2}$ for minimum mass solar nebula (MMSN), we can estimate the migration rate for a Jupiter mass planet as $\dot{r}_{\rm p}/r_{\rm p}={2T}/{M_{\rm p}\Omega_{\rm p}r_{\rm p}^{2}}\simeq 3.8\times 10^{-5}\Omega_{\rm p}$, which corresponds to a migration timescale of $\simeq 4200\ {\rm yrs}$. The migration timescale for the super-Earth is also on the same order of magnitude. Such a short timescale suggests an efficient process for the trapping of the planets within the rim.
@@
This work is partly supported by the National Natural Science Foundation of China (Grant No. 12133005, Z.Y.C. and S.M.). Y.P.L is supported in part by the National Natural Science Foundation of China (Grants NO. 12373070, and 12192223), the National Natural Science Foundation of Shanghai (Grant NO. 23ZR1473700). The authors acknowledge the High-Performance Computing platform at Westlake University for providing computational and data storage resources that have contributed to the results in this paper. The calculations have made use of the High Performance Computing Resource in the Core Facility for Advanced Research Computing at Shanghai Astronomical Observatory.

### v3

@@ III.1 Jupiter Migration
There is a general trend that the planet would migrate outward when the rim is located at $\Delta r_{\rm dz\_out}>-\Delta r_{\rm dz\_in}$, while the planet would migrate inward if the rim location satisfies $\Delta r_{\rm dz\_out}<-\Delta r_{\rm dz\_in}$ (see Panel b in Fig. 1). There is nearly zero migration torque when the planet is located at the rim center where $\Delta r_{\rm dz\_out}\simeq-\Delta r_{\rm dz\_in}$, shown as the dotted line in the panel c of Fig. 1. This simply suggests the dead zone center ($r_{\rm mid}=(r_{\rm dz\_in}+r_{\rm dz\_out})/2$) as the critical boundary of the migration torque reversal. Spatial gradients in the color map reveal smooth variations in the migration torque in terms of both its amplitude and sign. The size of the symbol reveals the variance of the torque, which could be related to the vortex-induced torque variations, as we show below.
@@
III.2 Super-Earth Migration
@@
For super-Earths, we can apply semi-analytical calculations by the formula for type I migration (Section II.2.2). The results for the comparison between theoretical and simulated torque values of the models enclosed by the green rectangle in Fig. 1(d) are illustrated in Fig. 4.
Figure 4: Theoretical calculation of the torque for super-Earth. Here, we set $M_{\rm p}=10M_{\oplus}$. We compute the theoretical torque values for the models enclosed by the green rectangle in Fig. 1(d). All these models have a dead zone width of $0.7r_{0}$, and the middle line of the surface density bump $r_{\rm dz\_mid}$ is shown on the horizontal axis of the figure. The blue diamond-dashed line represents the Lindblad torque, and the red square-dashed line denotes the corotation torque. The black dot-solid line shows the sum of the two obtained from analytical calculation, representing the total torque. The purple x-solid line corresponds to the total torque directly computed from Eq. 6.
@@
Figure 5: Theoretical calculations of the torque on a super-Earth for different disk surface density profiles. Here, we set $M_{\rm p}=10M_{\oplus}$. The theoretical torque values are computed at $\Delta r_{\rm mid}/r_{0}=-0.15$ in Fig. 4. This model feature a dead zone with an inner boundary at $0.8r_{0}$ and an outer boundary at $1.5r_{0}$. Panel a illustrates the variation in the slope of the surface density profile (the power-law index) for different values of $\delta$ in Eq. 3. Panel b displays the corresponding theoretical and simulated torques. The blue diamond-dashed line represents the Lindblad torque, and the red square-dashed line denotes the corotation torque. The black dot-solid line shows the sum of the two from analytical calculation, representing the total torque. The purple x-solid line corresponds to the total torque computed directly from Eq. 6.
We acknowledge that corotation torques could be sensitive to the gas density gradient. Therefore, we vary the slope of the density transition in the model at $\Delta r_{\rm mid}/r_{0}=-0.15$ in Fig. 4. This model features a dead zone with an inner boundary at $0.8r_{0}$ and an outer boundary at $1.5r_{0}$. As shown in Fig. 5(a), adopting different values of $\delta$ in Eq. 3 leads to variations in the slopes of the surface density profile (i.e., the power-law index). The slope of the ascending part ranges from 1.37 to 2.91, while the slope of the descending part ranges from 4.79-4.79 to 6.76-6.76. Fig. 5(b) presents the theoretical calculations of the corotation torque and Lindblad torque, along with a comparison between the total theoretical torque and the simulation results. Although slight variations in the corotation torque are observed, it remains the dominant contribution to the total torque even when the surface density transition becomes shallower.
@@ III.3 Effect of Vortex
The formation of vortices in protoplanetary disks is primarily driven by instabilities that arise under specific disk conditions. Vortices can be generated through multiple mechanisms, including Rossby Wave Instability (RWI) at the edges of gaps carved by massive planets in low-viscosity disks, at the boundaries of accretionally inactive dead zones, via the influence of binary companions, or through baroclinic instability. The RWI (Lovelace et al., 1999; Li et al., 2000, 2001; Lyra and Lin, 2013) is a dominant mechanism responsible for generating the large-scale, dust-trapping vortices observed in transitional disks. The RWI is characterized as an “edge mode” instability, analogous to the Kelvin-Helmholtz instability, which operates at sharp pressure maxima within the disk. It efficiently converts excess shear into vorticity, leading to the formation of long-lived, coherent vortical structures (Balmforth and Korycansky, 2001; Koller et al., 2003; Li et al., 2005, 2020).
In this study, vortex formation results from the combined effects of Jupiter’s gap-opening process and the ring-shaped surface density bump. The gap carved by Jupiter produces an annular region of reduced surface density, while the annular surface density bump is an accumulation of mass. The relative position and interaction between these two features determine both the occurrence and strength of any resulting vortices. For vortices at dead zone edges is primarily driven by gas accumulates at this viscosity transition, to maintain a radially constant accretion rate (${\dot{M}}\sim\nu\Sigma_{g}$), leading to the formation of a local maximum in the gas pressure profile (a “bump”). When this pressure bump is sufficiently sharp (with a width $\Delta_{\rm DZ}\lesssim 2H$, where $H$ is the local scale height), it becomes unstable to the RWI, which non-linearly evolves into long-lived, anticyclonic vortices (Miranda et al., 2017). This vortex generation mechanism at the dead zone edge is self-sustaining because the ongoing accretion flow continuously replenishes the gas at the bump, allowing the RWI and the vortices to persist for thousands of orbits.
@@
III.4 Effect of Planet Accretion
@@
According to Li et al. (2024), when the disk’s $\alpha$ parameter exceeds 0.003, accretion causes the planet to transition from inward to outward migration. In our setup, the dead zone has an $\alpha$ of $10^{-4}$, while the active region has an $\alpha$ of $10^{-2}$. The $\alpha$ in the active region exceeds the critical value required for accretion to exert an influence, whereas the $\alpha$ in the dead zone is insufficient for accretion to have a significant effect. Although accretion enhances the outward migration tendency of Jupiter, it is not sufficient to reverse the overall migration direction.
@@
III.5 The Cases with a Surface Density Gap
@@
Figure 8: The migration behavior of planets in a disk with a surface density gap. We inverted the $\alpha$ profile by swapping the values in the originally defined dead and active zones, while maintaining a smooth transition between them. The top row of panels (panel a–c) displays the migration map for Jupiter, whereas the bottom row (panel d–f) corresponds to super-Earths. For the upper column (left to right): the panel a employs $r_{\mathrm{dz\_in}}=0.8\,r_{0}$, $r_{\mathrm{dz\_out}}=1.4\,r_{0}$; the panel b uses $r_{\mathrm{dz\_in}}=0.7\,r_{0}$, $r_{\mathrm{dz\_out}}=1.3\,r_{0}$; and the panel c is specified as $r_{\mathrm{dz\_in}}=0.6\,r_{0}$, $r_{\mathrm{dz\_out}}=1.2\,r_{0}$. For the lower column (left to right): the panel d adopts $r_{\mathrm{dz\_in}}=1.0\,r_{0}$, $r_{\mathrm{dz\_out}}=1.6\,r_{0}$; the panel e has $r_{\mathrm{dz\_in}}=0.7\,r_{0}$, $r_{\mathrm{dz\_out}}=1.3\,r_{0}$; and the panel f corresponds to $r_{\mathrm{dz\_in}}=0.4\,r_{0}$, $r_{\mathrm{dz\_out}}=1.0\,r_{0}$. The selected configurations, indicating the relative positions between the surface density gap and the planet, are marked by diamond symbols in Fig. 1(c) and (d). The resulting migration directions, as shown in the left panel, are opposite to those shown in Fig. 1(b).
As shown in Fig. 8, panels a–c present the migration behavior of Jupiter in a disk with an annular surface density gap, while panels d–f show the corresponding results for super-Earths. The relative positions of the planets with respect to the surface density gap in these configurations are marked by diamond symbols in Fig. 1(c) and (d). For Jupiter, when located within the ring-shaped surface density gap, its migration direction is opposite to that observed in the annular surface density bump: with respect to the center of the surface density gap, planets migrate outward in the interior region and inward in the exterior region, causing Jupiter to converge toward the gap center. Moreover, gap opening by Jupiter further deepens and reinforces the pre-existing intrinsic surface density gap. For super-Earths, planets inside the ring-shaped surface density gap predominantly undergo inward migration. However, on or slightly outside the outer boundary of the surface density gap, outward migration occurs, whereas inward migration persists near the inner boundary. This implies that super-Earths are unlikely to remain trapped within the ring-shaped surface density gap, suggesting that intrinsic surface density gaps may not harbor super-Earths.
@@
III.6 Implementation of Artificial Azimuthal Torque
Recent advances in migration theory demonstrate that migration behavior is dependent on the $\alpha$ viscosity parameter (Paardekooper et al., 2023). The use of an $\alpha$ dead zone to mimic surface density bumps and gaps carries the risk of neglecting the potential effects of vortices on planetary migration. Therefore, we also employed an alternative method as a supplementary approach, applying an artificial torque to the protoplanetary disk to generate a surface density bump instead of an $\alpha$ dead zone.
To investigate the disk’s response to localized angular momentum redistribution, we implement an artificial torque function $T_{\rm art}(r)$ centered at a fiducial radius $r_{\rm mid}$. The torque profile is designed to converge material toward $r_{\rm mid}$, creating a density bump (or gap, depending on the sign) at this location.
The artificial torque is defined as:
$T_{\rm art}(r)=\begin{cases}-T_{0}\sin(x)\times 10^{-5},&\text{if }|r-r_{\rm mid}|\leq W\\ 0,&\text{otherwise}\end{cases}$ (13)
where $T_{0}$ is the torque amplitude, $W$ is the effective width of the torque application region, and the normalized radial coordinate $x$ is given by:
$x=\frac{r-r_{\rm mid}}{W}\pi$ (14)
The sinusoidal form ensures that the torque smoothly vanishes at the boundaries ($|r-r_{\rm mid}|=W$) while making the surface density profile reaching maximum strength at $r=r_{\rm mid}$. The negative sign in the expression ensures a converging torque profile: for $r<r_{\rm mid}$ (where $x<0$), $\sin(x)>0$, producing a positive (outward-directed) torque that pushes material outward toward $r_{\rm mid}$; conversely, for $r>r_{\rm mid}$ (where $x>0$), $\sin(x)<0$, producing a negative (inward-directed) torque that pulls material inward toward $r_{\rm mid}$.
Figure 9: The surface density profile derived from the artificial torque method. The solid line represents the surface density, and circular markers indicate the planet’s location. The color bar shows the magnitude of the average torque exerted on the planet for the given surface density profile and planetary position. The insets in the lower-left and lower-right corner, which zoom in on the region enclosed by the purple box, illustrate the direction of planetary migration. Panel a illustrates the Jupiter case, in which the planet migrates into the surface-density gap. Panel b shows the super-Earth case, where the planet moves toward the surface-density bump.
We examine four representative cases for Jupiter migration around a surface density gap. The Jupiter is placed at $r=1\ r_{0}$, with a gap width $W=0.5\ r_{0}$, amplitude $T_{0}=5\Sigma_{0}r^{4}_{0}\Omega^{2}_{0}$, and mid-line locations $r_{\mathrm{mid}}$ set to $1.4r_{0}$, $1.3r_{0}$, $1.2r_{0}$, and $1.1r_{0}$, respectively. Each configuration is then rescaled so that its mid-line coincides with $r=r_{0}$. Consequently, the planet positions in these four cases become $r_{\mathrm{p}}=0.91r_{0}$, $r_{\mathrm{p}}=0.83r_{0}$, $r_{\mathrm{p}}=0.77r_{0}$, and $r_{\mathrm{p}}=0.71r_{0}$, as illustrated in Fig. 9(a). The solid line traces the surface-density profile, and a circular marker indicates the planet location. The color bar encodes the magnitude of the time-averaged torque experienced by the planet for the given density structure and orbital radius. An inset in the lower-left corner, zooming into the region outlined by the purple box, illustrates the direction of planetary migration. From the computed average torques in these four cases, we confirm that Jupiter indeed converge toward the surface density gap, as shown by the migration directions in the inset of Fig. 9(a).
We also examine four representative cases for super-Earth migration around a surface density bump. The super-Earth is placed at $r=1\ r_{0}$, with a bump width $W=1\ r_{0}$, amplitude $T_{0}=10\Sigma_{0}r^{4}_{0}\Omega^{2}_{0}$, and mid-line locations $r_{\mathrm{mid}}$ set to $0.9r_{0}$, $0r_{0}, 1.1$$1r_{0}$, and $1.2r_{0}$, respectively. Each configuration is then rescaled so that its mid-line coincides with $r=r_{0}$. Consequently, the planet positions in these four cases become $r_{\mathrm{p}}=1.1r_{0}$, $r_{\mathrm{p}}=1.0r_{0}$, $r_{\mathrm{p}}=0.9r_{0}$, and $r_{\mathrm{p}}=0.8r_{0}$, as illustrated in Fig. 9(b). We also concluded that super-Earth converge toward the surface density gap.
IV Conclusion and Discussions
@@
Our calculations are scale-free; however, we can provide an example application to observational estimates. For instance, the cumulative torque in the top-left panel of Fig. 3 is approximately $T\sim 10^{-4}\Sigma_{0}r_{0}^{4}\Omega_{0}^{2}$. We adopted typical ALMA-observed locations and densities of rings to estimate the migration timescale (Andrews et al., 2018; Yoshida et al., 2025). By adopting an orbital radius of $r_{0}=50\,\text{AU}=50\times 1.496\times 10^{13}\,\text{cm}$, a surface density of $\Sigma_{0}=20\,\text{g/cm}^{2}$, we can estimate the migration rate for a Jupiter mass planet as $\dot{r}_{\rm p}/r_{\rm p}={2T}/{M_{\rm p}\Omega_{\rm p}r_{\rm p}^{2}}\simeq 1.1\times 10^{-3}\Omega_{\rm p}$, which corresponds to a migration timescale of $\simeq 5\times 10^{4}\ {\rm yrs}$. The migration timescale for the super-Earth is also on the same order of magnitude. Such a short timescale suggests an efficient process for the trapping of the planets within the rim.
@@
We thank the anonymous referee for their valuable suggestions and comments, which have significantly improved the manuscript. This work is partly supported by the National Natural Science Foundation of China (Grant No. 12133005, Z.Y.C. and S.M.). Y.P.L is supported in part by the National Natural Science Foundation of China (Grants NO. 12373070, and 12192223), the National Natural Science Foundation of Shanghai (Grant NO. 23ZR1473700). The authors acknowledge the High-Performance Computing platform at Westlake University for providing computational and data storage resources that have contributed to the results in this paper. The calculations have made use of the High Performance Computing Resource in the Core Facility for Advanced Research Computing at Shanghai Astronomical Observatory.